What is the vapor pressure of an aqueous solution that has a solute mole fraction of 0.300? The vapor pressure of water is 25.756 mmHg at 25 °C.


1
Expert's answer
2021-04-27T07:49:32-0400

χsolvent = 1.0000 - 0.3000 = 0.7000

Use Raoult's Law:

Psolution = (χsolvent) (P°solvent)

x = (0.7000) (25.756)

x = 18.03 mmHg